OpenStudy (anonymous):

Find the equation of a line through (4, -3) that is parallel to the line that passes through these points (-4,7) and (2,4). Write your answer in standard form (Ax + By = C). Show all your work. I'm really stuckkkk

OpenStudy (ddcamp):

To find the slope between the 2 points (x₁,y₁) and (x₂,y₂), use: \[m = \frac{ y₂ - y₁ }{ x₂ - x₁ }\] If two lines are parallel, their slopes are the same.

OpenStudy (anonymous):

But I need to find the equation of the line..

OpenStudy (ddcamp):

Once you know the slope, you can use the point-slope formula to find the equation: \[(y-y₁) = m(x-x₁)\]

OpenStudy (ddcamp):

Then you can solve it for standard form

OpenStudy (anonymous):

so would my formula to find the slope be (4-7)/2- -4)

OpenStudy (ddcamp):

Yup, that would be the slope (m).

OpenStudy (anonymous):

My answer comes up with -1/2 for the slope is that correct?

OpenStudy (ddcamp):

Yes, that is correct.

OpenStudy (anonymous):

What do I do next? Sorry I'm just really stuck

OpenStudy (ddcamp):

Now that you have the slope, plug it into the equation (y-y₁) = m(x-x₁) Where (x₁,y₁) is any point on the new line.

OpenStudy (anonymous):

Wait how do I plug it into the equation?

OpenStudy (ddcamp):

Put -1/2 in place of m, 4 in place of x₁, and -3 in place of y₁.

OpenStudy (anonymous):

So it would now be y- -3 = -1/2 (x-4)?

OpenStudy (ddcamp):

Yes.

OpenStudy (ddcamp):

Now put it into standard form by placing the x and y terms on the left, and the constants on the right.

OpenStudy (anonymous):

I'm stuck? what do you mean?

OpenStudy (anonymous):

(Ax + By = C). this is the example of standard form

OpenStudy (ddcamp):

y + 3 = -(1/2)x + 2 y + (1/2)x = -1 2y + x = -2

OpenStudy (anonymous):

THANK YOUUUUUUUU